OpenStudy (anonymous):

Use the following to determine the distance A (13,7) B (-9,4).

OpenStudy (saifoo.khan):

22.2

OpenStudy (anonymous):

The distance between 2 points is the hypotenuse of the right triangle whose sides are (Bx-Ax, By-Ay). The hypotenuse of a right triangle of sides (a,b) is \[c=sqrt(a^2+b^2)\]So in your case it's \[sqrt((-9-13)^2+(4-7)^2)\]\[=sqrt((-22)^2+(-3)^2)\]\[=22.2036033...\]